OpenStudy (anonymous):

Find the sum of the given arithmetic series.18 + 22 + 26 + 30 + ... + 50 = ?

OpenStudy (anonymous):

18+22=40 so that gives us 40+26+30+50 40+26=66 66+30=96 96+50= 146 Unless of course the ... between 30 and 50 is a missing number. Which means that is not the answer